OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 1 4 5 25

OpenStudy (compassionate):

Hello, my name is Stephen and I will be your instructor. The first step is distributing your terms. Do you see any terms that can be distributed? (hint, look at the terms in parenthesis)

OpenStudy (anonymous):

yes

OpenStudy (compassionate):

Can you distribute that for me, hun?

OpenStudy (anonymous):

3x+3=-2x+2+6

OpenStudy (compassionate):

Okay, now add like terms. (hint, the left side has NO like terms)

OpenStudy (anonymous):

x+3=8

OpenStudy (compassionate):

3x+3=-2x+2+6 Oops, you made a mistake! You want to add the 2 + 6 3x + 3 = -2x + 8 Now, you want to add the -2x 5x + 3 = 8

OpenStudy (anonymous):

its one